# Adaptive Learn-then-Test: Statistically Valid and Efficient Hyperparameter Selection

### v1

@@ Abstract
We introduce adaptive learn-then-test (aLTT), an efficient hyperparameter selection procedure that provides finite-sample statistical guarantees on the population risk of AI models. Unlike the existing learn-then-test (LTT) technique, which relies on conventional p-value-based multiple hypothesis testing (MHT), aLTT implements sequential data-dependent MHT with early termination by leveraging e-processes. As a result, aLTT can reduce the number of testing rounds, making it particularly well-suited for scenarios in which testing is costly or presents safety risks. Apart from maintaining statistical validity, in applications such as online policy selection for offline reinforcement learning and hyperparameter tuning for engineering systems, aLTT is shown to achieve the same performance as LTT while requiring only a fraction of the testing rounds.
@@ 1.1 Context and Motivation
Figure 1: Given a candidate pool $\Lambda$ of hyperparameters $\lambda$, which may be obtained through domain knowledge and/or hyperparameter optimization routines, aLTT applies a sequence of data-dependent testing rounds to select a subset $\hat{\Lambda}^{\rm rel}\subseteq\Lambda$ of hyperparameters that are deemed to be reliable. Each testing round $t$ evaluates through held-out data or real-world testing a subset of hyperparameters $\mathcal{I}^{t}\subseteq\Lambda$. An additional post-calibration selection step can be applied to choose a single hyperparameter $\hat{\lambda}$ in the selected subset $\hat{\Lambda}^{\rm rel}$ based on any performance criterion.
@@
The safe and reliable deployment of AI applications, or apps for short, hinges on the possibility of certifying their performance (Seshia et al., 2022; Tegmark and Omohundro, 2023). Depending on the problem, this may require controlling the missed detection probability in medical imaging (Lu et al., 2022; Mehrtash et al., 2020), ensuring safety measures for control policies (Lindemann et al., 2023; Zecchin et al., 2024), or verifying the correctness of the answers given by a large language model (Quach et al., 2023).
In practice, before deployment, AI apps can be often calibrated by selecting hyperparameters based on data set aside for this purpose or based on rounds of real-world testing. The hyperparameters can determine architectural choices, such as the composition of individual AI apps or the selection of submodules (Nikoloska and Simeone, 2022; Pfeiffer et al., 2023); inference parameters, such as the posterior temperature (Wenzel et al., 2020); optimization parameters, such as the learning rate for fine-tuning (Finn et al., 2017); implementation settings, such as the fidelity level of a simulator (Mungari et al., 2024); and post-processing parameters, such as the threshold used to set error bars (Shafer and Vovk, 2008).
@@
The learn-then-test (LTT) framework introduced by Angelopoulos et al. (2021) treats the calibration task of hyperparameter selection as a multiple hypothesis testing (MHT) problem. Accordingly, it associates each hyperparameter in a candidate set to the null hypothesis that the hyperparameter does not meet a reliability requirement on the population risk. Hypotheses are tested using p-values, and the probability of mistakenly detecting a hyperparameter as reliable is guaranteed via family-wise error rate (FWER)-controlling statistical procedures (Keselman and Rogan, 1977). This way, LTT ensures finite-sample, high-probability guarantees on the population risk of the selected hyperparameters.
@@
To prevent p-hacking (Head et al., 2015), LTT’s guarantees apply only to non-adaptive MHT procedures. However, related work on hyperparameter optimization has shown the significant benefits that can be accrued by adaptive exploration strategies that test hyperparameters sequentially in a data-driven manner (Swersky et al., 2014; Rakotoarison et al., 2024). This work aims at improving the data efficiency of LTT by leveraging recent advances in sequential MHT based on e-processes (Vovk and Wang, 2021; Waudby-Smith and Ramdas, 2024; Xu et al., 2021).
@@ 1.2 Related Work
Finite-sample statistical guarantees on inferential outputs can be obtained via conformal prediction methods (Shafer and Vovk, 2008; Angelopoulos et al., 2023) and, more generally, via conformal risk control and risk-controlling set-valued predictions (Angelopoulos et al., 2022; Bates et al., 2021). These methods calibrate classification or regression models by setting a threshold hyperparameter on the basis of held-out data to control the size of a prediction set. Calibrated predictors can be leveraged in predict-then-optimize control tasks to offer reliability guarantees (Vovk and Bendtsen, 2018; Lindemann et al., 2023; Zecchin et al., 2024).
@@
Beyond prediction sets, the problem of calibrating an AI app via the selection of hyperparameters from a candidate pool has been addressed through the LTT framework (Angelopoulos et al., 2021). As shown in Figure 1, the initial set of candidate hyperparameters can be obtained by using any hyperparameter optimization procedure, to be further discussed below. Leveraging p-value-based MHT via FWER-controlling procedures, LTT selects a subset of candidates that come with high-probability population risk guarantees.
@@
FWER guarantees are often too conservative, potentially resulting in empty calibration sets. The false discovery rate (FDR) is an alternative and less strict criterion that is often preferred for MHT in fields such as genetics (van den Oord and Sullivan, 2003), neuroimaging (Genovese et al., 2002), online advertising (Berman and Van den Bulte, 2022), and finance (Harvey and Liu, 2020).
@@
E-values have gained popularity in MHT due to their advantages over p-values (Vovk and Wang, 2021; Shafer and Vovk, 2019; Ramdas et al., 2020). Similarly to p-values, e-values measure the statistical plausibility of the null hypothesis. Specifically, an e-value can be thought of as a special type of p-value that has additional robustness properties. Specifically, the key property of interest in this paper is that, unlike p-values, e-values can be readily combined to obtain e-processes, making it possible to devise sequential testing strategies with anytime safety (Wang and Ramdas, 2022; Ramdas et al., 2023). E-processes have been applied to problems such as sequential change detection (Shin et al., 2022), multiple bandit testing (Xu et al., 2021), two-sample testing (Shekhar and Ramdas, 2023), and mean estimation of bounded random variables (Waudby-Smith and Ramdas, 2024).
@@
Hyperparameter optimization is a vast field focused on the optimization of the hyperparameters of training algorithms, such as learning rate, weight decay, and dropout rate (Swersky et al., 2014; Pedregosa, 2016; Maclaurin et al., 2015; Lindauer et al., 2022). As seen in Figure 1, hyperparameter optimization typically operates in continuous domains, and it can serve as a preliminary step for the identification of candidate hyperparameters. While hyperparameter optimization does not provide statistical guarantees on the population risk, the goal of hyperparameter selection methods such as LTT is to formally test a subset of candidate hyperparameters for statistical validity.
@@ 1.3 Main Contributions
$\bullet$ We study two practical scenarios requiring hyperparameter selection, namely online policy selection for offline reinforcement learning (Fujimoto and Gu, 2021) and resource allocation for wireless engineering (Stanczak et al., 2009). In both cases, aLTT is shown to deliver reliable and effective hyperparameters using only a small fraction of the testing rounds required by LTT.
@@ 2.1 Setting
As illustrated in Figure 1, assume that a discrete subset $\Lambda=\{\lambda_{1},\dots,\lambda_{N}\}$ of hyperparameters has been identified based on domain knowledge and/or conventional hyperparameter optimization methods. For a user-specified reliability level $\alpha\in[0,1]$, we aim at determining the subset of hyperparameters in set $\Lambda$ that conforms with the required reliability level $\alpha$, i.e.,
@@ 2.2 Performance Criteria
Definition 1 ($(\alpha,\delta)$-FWER-controlling set).
@@
Definition 2 ($(\alpha,\delta)$-FDR-controlling set).
@@ 2.3 Sequential and Adaptive Hyperparameter Selection
LTT uses the accumulated evidence at the end of the testing process, $\mathcal{D}^{T}$, to compute a valid p-value for the null hypothesis (7) using, for instance, the Hoeffding-Bentkus concentration inequality introduced in Bates et al. (2021). Based on the collection of p-values $\mathcal{P}=\{P_{i}\}_{i=1}^{N}$, LTT selects a subset of hyperparameters $\hat{\Lambda}^{\rm LTT}$ using a FWER-controlling algorithm $\mathcal{A}^{\rm FWER}(\mathcal{P})$. A variant of LTT that is FDR-controlling can be readily obtained by using an FDR-controlling selection rule $\mathcal{A}^{\rm FDR}(\mathcal{P})$. Examples of FWER and FDR controlling procedures are provided in the Supplementary Material.
@@
Candidate hyperparameters $\Lambda$, prior knowledge $\mathcal{D}^{0}$, reliability level $\alpha$, error tolerance level $\delta$, acquisition policy $\{\mathcal{Q}^{t}\}_{t\geq 1}$, FWER/FDR-controlling selection rule $\mathcal{A}^{\rm FWER}$/$\mathcal{A}^{\rm FDR}$, betting strategy $\{\mu^{t}_{i}\}_{t\geq 1,i=1,\dots,N}$, maximum number of iterations $t_{\rm max}$ and minimal hyperparameter set cardinality $d$
$(\alpha,\delta)$-FWER/FDR-controlling hyperparameter set $\hat{\Lambda}^{{\rm aLTT},T}$
@@
     $t\leftarrow t+1$ return $\hat{\Lambda}^{{\rm aLTT},T}$
@@ 4.1 Hypothesis Testing via E-Processes
$\Pr\left[\frac{1}{E_{i}}\leq x\bigg{|}\mathcal{H}_{i}\right]\leq\mathbb{E}[E_{i}|\mathcal{H}_{i}]x\leq x.$ (10)
For each null hypothesis $\mathcal{H}_{i}$, given an observation $Z$ and a fixed $\mu\in(0,1/(1-\alpha))$, a valid e-value is given by (Waudby-Smith and Ramdas, 2024)
@@
where the betting strategy $\mu_{i}^{t}\in(0,1/(1-\alpha))$ can be optimized as a function of the past risk estimates $\{R(\lambda_{i},Z^{\tau}_{i})\}_{\tau<t}$ and $E^{0}_{i}=1$. Based on the discussion above, the e-process (12) represents the wealth accumulated up to time $t$ by a gambler making sequential bets $\{\mu^{\tau}_{i}\}_{\tau<t}$ (Shafer and Vovk, 2019; Waudby-Smith and Ramdas, 2024). As such, the gambler’s wealth up to time $t$, $E^{t}_{i}$, provides evidence against the null hypothesis that the hyperparameter $\lambda_{i}$ is unreliable.
@@ 4.2 Adaptive Acquisition Policy
With this information, a prediction set $\hat{\Lambda}^{{\rm aLTT},t}$ is evaluated by employing either an FWER-controlling method $\mathcal{A}^{\rm FWER}(\mathcal{P}^{t})$ based on the p-values $\mathcal{P}^{t}=\{P^{t}_{i}\}_{i=1}^{N}$ in (13); or an FDR-controlling procedure $\mathcal{A}^{\rm FDR}(\mathcal{E}^{t})$, such as the e-Benjamini-Hochberg (eBH) method, reviewed in the Supplementary Material (Wang and Ramdas, 2022), using directly the e-values (14).
@@ 5.1 Online Policy Selection for Offline Reinforcement Learning
Offline reinforcement learning enables the training of control policies based on a fixed data set collected by using a possibly unknown behavior policy, without any online interaction with the environment (Levine et al., 2020). However, the estimate of the performance of the trained policies obtained from offline data can differ substantially from the actual performance in the real world. This makes it practically essential to validate the policies’ performance via online interaction with the environment (Liu et al., 2023).
To reduce the cost and potential harm of online validation procedures, the number of online interactions of the pre-trained candidate policies with the real world must be kept to a minimum (Garcıa and Fernández, 2015). To this end, in this subsection, we investigate the potential benefits of the proposed aLTT framework as a means to select a subset of candidate policies that enjoy performance guarantees with respect to the real-world environment.
@@ 5.1.2 Results
In our experiments, we consider the Half Cheetah control problem from the OpenAI Gym MuJoCo tasks (Todorov et al., 2012) and use control policies obtained via the offline reinforcement learning algorithm TD3+BC (Fujimoto and Gu, 2021). The TD3+BC algorithm leverages an offline dataset $\mathcal{D}$ to optimize policies by maximizing the standard deterministic policy gradient objective (Silver et al., 2014), combined with a behavioral cloning regularization term, whose strength is controlled by a hyperparameter $\lambda$ in (Silver et al., 2014, Eq. 5). We produce $N=20$ different control policies by setting the hyperparameter $\lambda$ in the TD3+BC training objective on an evenly spaced grid in the interval $[0.25,5]$. Unless stated otherwise, we consider a target reliability $\alpha=0.57$ and a target FDR requirement $\delta=0.1$.
@@
Finally, the value of the parameter $\mu^{t}_{i}$ in aLTT is set by following the approximate growth rate adaptive to the particular alternative (aGRAPA) betting strategy in Waudby-Smith and Ramdas (2024) with other adaptive and non-adaptive betting strategies evaluated in the Supplementary Material.
@@
5.2 Reliable Hyperparameter Selection for Wireless Resource Allocation

### v2

@@ Abstract
We introduce adaptive learn-then-test (aLTT), an efficient hyperparameter selection procedure that provides finite-sample statistical guarantees on the population risk of AI models. Unlike the existing learn-then-test (LTT) technique, which relies on conventional p-value-based multiple hypothesis testing (MHT), aLTT implements sequential data-dependent MHT with early termination by leveraging e-processes. As a result, aLTT can reduce the number of testing rounds, making it particularly well-suited for scenarios in which testing is costly or presents safety risks. Apart from maintaining statistical validity, in applications such as online policy selection for offline reinforcement learning and prompt engineering, aLTT is shown to achieve the same performance as LTT while requiring only a fraction of the testing rounds.
Machine Learning, ICML
@@ 1.1 Context and Motivation
Figure 1: An example application of aLTT to reliable prompt optimization (Zhou et al., 2023; Quach et al., 2023; Schneider et al., 2024). A set $\Lambda$ of candidate prompts for a movie recommender is generated using an LLM and/or prior experience. Prompts serve as an example of a discrete set of hyperparameters to be optimized using aLTT. The goal is to select a subset $\hat{\Lambda}^{\rm rel}\subseteq\Lambda$ of prompts that guarantee a sufficiently high recommendation accuracy. To this end, aLTT applies a sequence of data-dependent testing rounds with adaptive termination. Specifically, at each testing round $t$, aLTT estimates the performance of a subset of hyperparameters $\mathcal{I}^{t}\subseteq\Lambda$ through held-out data or real-world testing. The subset to be tested is selected based on prior testing outcomes, and the process stops as soon as a sufficiently large reliable subset $\hat{\Lambda}^{\rm rel}$ is identified. An additional post-calibration selection step can be applied to choose a single hyperparameter $\hat{\lambda}$ from the selected subset $\hat{\Lambda}^{\rm rel}$ based on users preferences.
@@
The safe and reliable deployment of AI applications, or apps for short, hinges on the possibility of certifying their performance (Seshia et al., 2022; Tegmark & Omohundro, 2023). Depending on the problem, this may require controlling the missed detection probability in medical imaging (Lu et al., 2022; Mehrtash et al., 2020), ensuring safety measures for control policies (Lindemann et al., 2023; Zecchin et al., 2024), or verifying the correctness of the answers given by a large language model (LLM) (Quach et al., 2023).
In practice, before deployment, AI apps can be often calibrated by selecting hyperparameters based on data set aside for this purpose or based on rounds of real-world testing. In the current scaling-centric era of large, Internet-scale, data sets (Xiao, 2024), hyperparameter optimization is typically modeled as a bandit problem in which the goal is the minimization of the training loss over a set of discrete configurations. The configurations may include prompt designs for the fine-tuning of large language models (LLMs) (Zhou et al., 2023; Quach et al., 2023; Schneider et al., 2024), architectural choices related to model scale (Nasir et al., 2024) or to the timing of early decisions (Schuster et al., 2022), or policy settings for offline reinforcement learning (Paine et al., 2020; Fujimoto & Gu, 2021).
@@
The learn-then-test (LTT) method introduced by (Angelopoulos et al., 2021) has recently emerged as a formal framework to address hyperparameter selection over a discrete space. LTT treats the calibration task of hyperparameter selection as a multiple hypothesis testing (MHT) problem. Accordingly, it associates each hyperparameter in a candidate set to the null hypothesis that the hyperparameter does not meet a reliability requirement on the population risk. Hypotheses are tested using p-values, and the probability of mistakenly detecting a hyperparameter as reliable is guaranteed via family-wise error rate (FWER)-controlling statistical procedures (Keselman & Rogan, 1977). This way, LTT ensures finite-sample, high-probability guarantees on the population risk of the selected hyperparameters.
@@
To prevent p-hacking (Head et al., 2015), LTT’s guarantees apply only to non-adaptive MHT procedures. However, related work on hyperparameter optimization has shown the significant benefits that can be accrued by adaptive exploration strategies that test hyperparameters sequentially in a data-driven manner (Swersky et al., 2014; Rakotoarison et al., 2024). This work aims at improving the data efficiency of LTT by leveraging recent advances in sequential MHT based on e-processes (Vovk & Wang, 2021; Waudby-Smith & Ramdas, 2024; Xu et al., 2021).
@@ 1.2 Related Work
Finite-sample statistical guarantees on inferential outputs can be obtained via conformal prediction methods (Shafer & Vovk, 2008; Angelopoulos et al., 2023) and, more generally, via conformal risk control and risk-controlling set-valued predictions (Angelopoulos et al., 2022; Bates et al., 2021). These methods calibrate classification or regression models by setting a threshold hyperparameter based on held-out data to control the size of a prediction set. Calibrated predictors can be leveraged in predict-then-optimize control tasks to offer reliability guarantees (Vovk & Bendtsen, 2018; Lindemann et al., 2023; Zecchin et al., 2024).
@@
Beyond prediction sets, the problem of calibrating an AI app via the selection of hyperparameters from a candidate pool has been addressed through the LTT framework (Angelopoulos et al., 2021). For example, as illustrated in Figure 1, for the problem of prompt engineering (Zhou et al., 2023), the initial set of candidate hyperparameters encompasses instruction prompt templates generated by an LLM and/or via prior experience. Leveraging p-value-based MHT via FWER-controlling procedures, LTT selects a subset of candidates that come with high-probability population risk guarantees.
@@
FWER guarantees are often too conservative, potentially resulting in empty calibration sets. The false discovery rate (FDR) is an alternative and less strict criterion that is often preferred for MHT in fields such as genetics (van den Oord & Sullivan, 2003), neuroimaging (Genovese et al., 2002), online advertising (Berman & Van den Bulte, 2022), and finance (Harvey & Liu, 2020).
@@
E-values have gained popularity in MHT due to their advantages over p-values (Vovk & Wang, 2021; Shafer & Vovk, 2019; Ramdas et al., 2020). Similarly to p-values, e-values measure the statistical plausibility of the null hypothesis. Specifically, an e-value can be thought of as a special type of p-value that has additional robustness properties. Specifically, the key property of interest in this paper is that, unlike p-values, e-values can be readily combined to obtain e-processes, making it possible to devise sequential testing strategies with anytime safety (Wang & Ramdas, 2022; Ramdas et al., 2023). E-processes have been applied to problems such as sequential change detection (Shin et al., 2022), multiple bandit testing (Xu et al., 2021), two-sample testing (Shekhar & Ramdas, 2023), and mean estimation of bounded random variables (Waudby-Smith & Ramdas, 2024).
@@
Hyperparameter optimization is a vast field focused on the optimization of the hyperparameters of training algorithms, such as learning rate, weight decay, and dropout rate (Swersky et al., 2014; Pedregosa, 2016; Maclaurin et al., 2015; Lindauer et al., 2022). Hyperparameter optimization typically operates in continuous domains, and it can serve as a preliminary step for the identification of candidate hyperparameters. While hyperparameter optimization does not provide statistical guarantees on the population risk, the goal of hyperparameter selection methods such as LTT is to formally test a subset of candidate hyperparameters for statistical validity.
@@ 1.3 Main Contributions
$\bullet$ We study two practical scenarios requiring hyperparameter selection, namely online policy selection for offline reinforcement learning (Fujimoto & Gu, 2021) and automated prompt engineering (Zhou et al., 2023). In both cases, aLTT is shown to deliver reliable and effective hyperparameters using only a small fraction of the testing rounds required by LTT.
@@ 2.1 Setting
As illustrated in Figure 1, for a given discrete subset $\Lambda=\{\lambda_{1},\dots,\lambda_{N}\}$ of hyperparameters and a user-specified reliability level $\alpha\in[0,1]$, we aim at determining the subset of hyperparameters in set $\Lambda$ that conforms with the required reliability level $\alpha$, i.e.,
@@ 2.2 Performance Criteria
Definition 2.1 ($(\alpha,\delta)$-FWER-controlling set).
@@
Definition 2.2 ($(\alpha,\delta)$-FDR-controlling set).
@@ 2.3 Sequential and Adaptive Hyperparameter Selection
LTT uses the accumulated evidence at the end of the testing process, $\mathcal{D}^{T}$, to compute a valid p-value for the null hypothesis (7) using, for instance, the Hoeffding-Bentkus concentration inequality introduced in (Bates et al., 2021). Based on the collection of p-values $\mathcal{P}=\{P_{i}\}_{i=1}^{N}$, LTT selects a subset of hyperparameters $\hat{\Lambda}^{\rm LTT}$ using a FWER-controlling algorithm $\mathcal{A}^{\rm FWER}(\mathcal{P})$. A variant of LTT that is FDR-controlling can be readily obtained by using an FDR-controlling selection rule $\mathcal{A}^{\rm FDR}(\mathcal{P})$. Examples of FWER and FDR controlling procedures are provided in the Supplementary Material.
@@
  Require: Candidate hyperparameters $\Lambda$, prior knowledge $\mathcal{D}^{0}$, reliability level $\alpha$, error tolerance level $\delta$, acquisition policy $\{\mathcal{Q}^{t}\}_{t\geq 1}$, FWER/FDR-controlling selection rule $\mathcal{A}^{\rm FWER}$/$\mathcal{A}^{\rm FDR}$, betting strategy $\{\mu^{t}_{i}\}_{t\geq 1,i=1,\dots,N}$, maximum number of iterations $t_{\rm max}$ and minimal hyperparameter set cardinality $d$
  Ensure: $(\alpha,\delta)$-FWER/FDR-controlling hyperparameter set $\hat{\Lambda}^{{\rm aLTT},T}$
@@
     end if
     $t\leftarrow t+1$
  end while
  Return: $\hat{\Lambda}^{{\rm aLTT},T}$
@@ 4.1 Hypothesis Testing via E-Processes
$\Pr\left[\frac{1}{E_{i}}\leq x\bigg{|}\mathcal{H}_{i}\right]\leq\mathbb{E}[E_{i}|\mathcal{H}_{i}]x\leq x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{ }\forall x\in[0,1].}$ (10)
For each null hypothesis $\mathcal{H}_{i}$, given an observation $Z$ and a fixed $\mu\in(0,1/(1-\alpha))$, a valid e-value is given by (Waudby-Smith & Ramdas, 2024)
@@
where the betting strategy $\mu_{i}^{t}\in(0,1/(1-\alpha))$ can be optimized as a function of the past risk estimates $\{R(\lambda_{i},Z^{\tau}_{i})\}_{\tau<t}$ and $E^{0}_{i}=1$. Based on the discussion above, the e-process (12) represents the wealth accumulated up to time $t$ by a gambler making sequential bets $\{\mu^{\tau}_{i}\}_{\tau<t}$ (Shafer & Vovk, 2019; Waudby-Smith & Ramdas, 2024). As such, the gambler’s wealth up to time $t$, $E^{t}_{i}$, provides evidence against the null hypothesis that the hyperparameter $\lambda_{i}$ is unreliable.
@@ 4.2 Adaptive Acquisition Policy
With this information, a prediction set $\hat{\Lambda}^{{\rm aLTT},t}$ is evaluated by employing either an FWER-controlling method $\mathcal{A}^{\rm FWER}(\mathcal{P}^{t})$ based on the p-values $\mathcal{P}^{t}=\{P^{t}_{i}\}_{i=1}^{N}$ in (13); or an FDR-controlling procedure $\mathcal{A}^{\rm FDR}(\mathcal{E}^{t})$, such as the e-Benjamini-Hochberg (eBH) method, reviewed in the Supplementary Material (Wang & Ramdas, 2022), using directly the e-values (14).
@@ 5.1 Online Policy Selection for Offline Reinforcement Learning
Offline reinforcement learning enables the training of control policies based on a fixed data set collected by using a possibly unknown behavior policy, without any online interaction with the environment (Levine et al., 2020). However, the estimate of the performance of the trained policies obtained from offline data can differ substantially from the actual performance in the real world. This makes it practically essential to validate the policies’ performance via online interaction with the environment (Paine et al., 2020; Liu et al., 2023).
To reduce the cost and potential harm of online validation procedures, the number of online interactions of the pre-trained candidate policies with the real world must be kept to a minimum (Garcıa & Fernández, 2015). To this end, in this subsection, we investigate the potential benefits of the proposed aLTT framework as a means to select a subset of candidate policies that enjoy performance guarantees with respect to the real-world environment.
@@ 5.1.2 Results
In our experiments, we consider the Half Cheetah control problem from the OpenAI Gym MuJoCo tasks (Todorov et al., 2012) and use control policies obtained via the offline reinforcement learning algorithm TD3+BC (Fujimoto & Gu, 2021). The TD3+BC algorithm leverages an offline data set $\mathcal{D}$ to optimize policies by maximizing the standard deterministic policy gradient objective (Silver et al., 2014), combined with a behavioral cloning regularization term, whose strength is controlled by a hyperparameter $\lambda$ in [Eq. 5](Silver et al., 2014). We produce $N=20$ different control policies by setting the hyperparameter $\lambda$ in the TD3+BC training objective on an evenly spaced grid in the interval $[0.25,5]$. Unless stated otherwise, we consider a target reliability $\alpha=0.57$ and a target FDR requirement $\delta=0.1$.
@@
Finally, the value of the parameter $\mu^{t}_{i}$ in aLTT is set by following the approximate growth rate adaptive to the particular alternative (aGRAPA) betting strategy in (Waudby-Smith & Ramdas, 2024) with other adaptive and non-adaptive betting strategies evaluated in the Supplementary Material.
@@
5.2 Reliable Automated Prompt Engineering
Prompt engineering focuses on designing and refining input instructions for LLMs (Reynolds & McDonell, 2021; Shin et al., 2020). Recent studies have demonstrated the effectiveness of methods that automate this search using LLMs as prompt generators (Zhou et al., 2023; Zhang et al., 2023). However, while LLMs are capable of producing high-quality instruction templates, at a level comparable to human annotators, supervision and testing remain essential to filter out poorly performing prompts.
In this section, we propose applying aLTT to automated prompt engineering (Zhou et al., 2023) to generate instructions with statistical performance guarantees. As illustrated in Figure 1, given a set $\Lambda$ of candidate instructions generated by an LLM, we use aLTT to sequentially test the instructions and identify a sufficiently large number of prompts that meet user-defined reliability requirements.